OpenStudy (anonymous):

cos80cos20+sin80sin20

jimthompson5910 (jim_thompson5910):

hint: cos(x-y) = cos(x)cos(y) + sin(x)sin(y)

OpenStudy (anonymous):

cos80+20

jimthompson5910 (jim_thompson5910):

let x = 80, y = 20

OpenStudy (anonymous):

then what

jimthompson5910 (jim_thompson5910):

cos(x-y) = cos(x)cos(y) + sin(x)sin(y) cos(x)cos(y) + sin(x)sin(y) = cos(x-y) cos(80)cos(20) + sin(80)sin(20) = cos(80-20) cos(80)cos(20) + sin(80)sin(20) = ???

OpenStudy (anonymous):

so cos 80+20

jimthompson5910 (jim_thompson5910):

what does cos(80-20) turn into

OpenStudy (anonymous):

cos60

jimthompson5910 (jim_thompson5910):

and what does cos(60) evaluate to?

OpenStudy (anonymous):

\[\sqrt{2}/\]

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

that's cos(45)

OpenStudy (anonymous):

1/2

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

cos(80)cos(20) + sin(80)sin(20) = cos(60) cos(80)cos(20) + sin(80)sin(20) = 1/2

OpenStudy (anonymous):

since i have you here can i ask you how to do \[\cos(2\sin^{-1} (-3\div5)\]

OpenStudy (anonymous):

i got -1/2 as the answer was i doing something right or just way off

jimthompson5910 (jim_thompson5910):

-1/2 is incorrect

OpenStudy (anonymous):

what do i do

OpenStudy (anonymous):

after the identity what is the next step

OpenStudy (anonymous):

i did so can i just swap the sin with the cos fractions

OpenStudy (anonymous):

4

OpenStudy (anonymous):

so the answer would be 4/5

jimthompson5910 (jim_thompson5910):

idk why but I've been doing this all wrong...

jimthompson5910 (jim_thompson5910):

let me fix it

jimthompson5910 (jim_thompson5910):

cos(2*arcsin(-3/5)) cos(2x) 2*cos^2(x) - 1 2*(4/5)^2 - 1 2*(16/25) - 1 32/25 - 1 32/25 - 25/25 7/25

jimthompson5910 (jim_thompson5910):

So, cos(2*arcsin(-3/5)) = 7/25

OpenStudy (anonymous):

thank you the 25/25 was pi correct?

OpenStudy (anonymous):

o sorry it was the 1

jimthompson5910 (jim_thompson5910):

pi?